### v2

# A Simple yet Powerful Instance-Aware Prompting Framework for Training-free Camouflaged Object Segmentation

Abstract
Camouflaged Object Segmentation (COS) remains highly challenging due to the intrinsic visual similarity between target objects and their surroundings. While training-based COS methods achieve good performance, their performance degrades rapidly with increased annotation sparsity. To circumvent this limitation, recent studies have explored training-free COS methods, leveraging the Segment Anything Model (SAM) by automatically generating visual prompts from a single task-generic prompt (e.g., ”camouflaged animal”) uniformly applied across all test images. However, these methods typically produce only semantic-level visual prompts, causing SAM to output coarse semantic masks and thus failing to handle scenarios with multiple discrete camouflaged instances effectively. To address this critical limitation, we propose a simple yet powerful Instance-Aware Prompting Framework (IAPF), the first training-free COS pipeline that explicitly converts a task-generic prompt into fine-grained instance masks. Specifically, the IAPF comprises three steps: (1) Text Prompt Generator, utilizing task-generic queries to prompt a Multimodal Large Language Model (MLLM) for generating image-specific foreground and background tags; (2) Instance Mask Generator, leveraging Grounding DINO to produce precise instance-level bounding box prompts, alongside the proposed Single-Foreground Multi-Background Prompting strategy to sample region-constrained point prompts within each box, enabling SAM to yield a candidate instance mask; (3) Self-consistency Instance Mask Voting, which selects the final COS prediction by identifying the candidate mask most consistent across multiple candidate instance masks. Extensive evaluations on standard COS benchmarks demonstrate that the proposed IAPF significantly surpasses existing state-of-the-art training-free COS methods.
Codehttps://github.com/ycyinchao/IAPF
Introduction
Figure 1: Motivation of the proposed IAPF. (a) Visualization of box prompts generated by MLLM, MaxIOUBox, and our method. Only the instance-aware strategy yields multiple instance-level boxes. (b) Quantitative results of the COD10K dataset showing our superior box prompts accuracy, especially under a high IoU threshold.
Camouflaged Object Segmentation (COS) holds significant research value and broad application prospects, including agriculture, medicine, and military reconnaissance. It remains a challenging yet essential task in computer vision, focusing on accurately identifying and segmenting objects visually indistinguishable from their surroundings. Traditional fully-supervised COS methods (Fan et al. 2020; Yin et al. 2024) heavily rely on dense pixel-level annotations, significantly escalating annotation costs and limiting practical applicability. To alleviate annotation burdens, weakly-supervised and unsupervised paradigms (Chen et al. 2024b; Du et al. 2025) have been introduced; however, these training-based methods typically suffer from performance degradation as label sparsity increases.
Recent advancements in promptable segmentation models (e.g., Segment Anything Model, SAM (Kirillov et al. 2023)) have demonstrated promising potential, achieving competitive results across various segmentation tasks with minimal manual visual prompts (i.e., points and boxes). Nevertheless, manual prompts significantly constrain their practical deployment when segmenting specific objects of interest. Consequently, increasing attention has been devoted to developing training-free segmentation methods that automatically generate visual prompts without additional supervision. An active research area within this context is training-free COS methods (Yin et al. 2025; Hu et al. 2024a, b; Tang et al. 2024), which utilizes a single task-generic prompt (e.g., ”camouflaged animal”) indiscriminately applied across all test samples within the COS domain to automatically generate visual prompts required by SAM, thus enabling segmentation of camouflaged objects of interest.
Despite such advancements, existing methods primarily produce semantic-level visual prompts rather than instance-level prompts, directly determining whether SAM outputs semantic or instance masks. According to SAM, explicit bounding boxes are required for instance segmentation. Thus, intuitively, generating instance-level bounding boxes is crucial. Unfortunately, current methods fail to meet this requirement adequately: 1) As illustrated in Fig. 1(a)(1), existing approaches employing Multimodal Large Language Models (MLLM) (Liu et al. 2024) via visual question-answering typically produce a single box covering all objects collectively. 2) Alternatively, methods using intermediate semantic masks (Fig. 1(a)(2)) adopt a MaxIOUBox approach, which identifies all connected components in the mask, computes bounding boxes, and retains the largest one. The above two approaches inevitably result in a single box per image, which typically covers only one or overlapping multiple instances, severely limiting their efficacy in scenarios containing multiple discrete camouflaged instances.
As demonstrated in Fig. 1(a)(3), our method leverages the instance-aware detection capability of Grounding DINO (Liu et al. 2023) to generate multiple instance-level bounding boxes. This enables SAM to perform fine-grained instance segmentation even in complex camouflage scenes. To quantify the accuracy of bounding box prompts generated by the above methods, we evaluate their performance in Fig. 1(b). It is evident that our approach significantly outperforms existing methods in bounding box prompts accuracy on the COD10K dataset (Fan et al. 2020), achieving a mAP of $0.662$ at a stringent IoU threshold of $0.75$, substantially surpassing the $0.009$ and $0.065$ achieved by the above two methods, respectively.
To overcome these limitations, we propose a simple yet powerful Instance-Aware Prompting Framework (IAPF) specifically designed for training-free Camouflaged Object Segmentation. As depicted in Fig. 2, the IAPF initially utilizes an MLLM to automatically generate sets of foreground and background category tags from the input image. Subsequently, leveraging these text prompts, the proposed Instance Mask Generator incorporates Grounding DINO (Liu et al. 2023) to accurately produce multiple instance-aware bounding box prompts, overcoming the critical shortcomings of existing methods. Additionally, within the Instance Mask Generator, we also introduce a region-constrained point prompting strategy named Single-Foreground Multi-Background Prompting (SFMBP), which efficiently leverages multiple background tags to further refine point prompts for instance segmentation. These carefully selected visual prompts (i.e., points and boxes) enable SAM to generate a precise and fine-grained candidate instance mask. Lastly, we introduce a Self-consistency Instance Mask Voting mechanism by evaluating multiple candidate instance masks, ensuring final robust and accurate COS results. Our primary contributions are summarized as follows:
  1. 1.
    We present a novel, training-free Instance-Aware Prompting Framework that effectively addresses the limitations of semantic-level visual prompts in existing training-free COS methods, enabling accurate and complete Camouflaged Object Segmentation.
  2. 2.
    We propose the Instance Mask Generator, an integral module that combines Grounding DINO and our SFMBP strategy to generate precise instance-level visual prompts, substantially improving the segmentation completeness and accuracy in complex camouflage scenarios.
  3. 3.
    Extensive experimental evaluations on multiple COS benchmarks demonstrate that the proposed IAPF significantly outperforms existing training-free COS methods.
Figure 2: Framework of the proposed IAPF, which consists of three steps: (1) Text Prompt Generator: an MLLM turns task-generic prompts and an image into foreground/background tags. (2) Instance Mask Generator (the lower part): Grounding DINO produces NN bbox prompts; the Single-Foreground Multi-Background Prompting (SFMBP) strategy uses CLIP to create foreground and multi-background heatmaps and samples region-constrained point prompts; SAM converts NN visual prompt pairs (i.e., points and boxes) to a candidate instance mask. (3) Self-consistency Instance Mask Voting mechanism: The instance mask with the highest consistency is voted as the final COS prediction.
Related Work
Camouflaged Object Segmentation
Early fully-supervised COS networks (Fan et al. 2020; Hu et al. 2023) achieve high accuracy by learning from dense pixel masks, but these masks are expensive to obtain, and every new domain requires additional annotation and retraining. Weakly-supervised alternatives (Chen et al. 2024a; He et al. 2023b; Li et al. 2025a) alleviate the burden by using sparse cues—-points, scribbles, or boxes—-yet their accuracy drops rapidly as the cues become scarce and they still need task-specific optimisation. Unsupervised pipelines (Zhang and Wu 2023; Shou et al. 2025; Du et al. 2025) remove labels altogether, but at the cost of a clear performance gap. In short, all existing COS families are training-centric, and their effectiveness is tightly coupled to the amount of supervision available, motivating a paradigm that can segment camouflaged objects without any training or annotation.
Segment Anything Model for COS
SAM (Kirillov et al. 2023) produces precise masks only when supplied with explicit visual prompts. Early COS adaptations fine-tune SAM with dense (fully-supervised (Chen et al. 2023; Gao et al. 2025)) or sparse (weakly-supervised (He et al. 2023a; Chen et al. 2024b)) labels, but these variants still require training. To remove this dependency, training-free pipelines automatically generate visual prompts for SAM at test time. GenSAM (Hu et al. 2024a) is the first to do so: it lets an MLLM produce text prompts, feeds them to CLIP to obtain visual prompts for SAM. ProMaC (Hu et al. 2024b) improves GenSAM by exploiting hallucination priors to reduce the inaccuracy of MLLM-generated text prompts. RDVP-MSD (Yin et al. 2025) targets the other source–noisy CLIP heatmaps–by constraining point sampling to a box, yet it still relies on a single semantic-level box prompt. Consequently, all these methods struggle when an image contains multiple camouflaged objects.
Method
Framework Overview
We propose the Instance-Aware Prompting Framework (IAPF), a training-free solution that segments multiple unseen classes of camouflaged objects from a single task-generic prompt. The key novelty is to supply instance-level visual prompts to SAM for predicting instance masks rather than coarse semantic masks. Given a test image $X$ and a task-generic prompt $p_{i}^{g}$ (e.g., ”camouflaged animal”), the IAPF proceeds in three frozen steps. (1) Text Prompt Generator. An MLLM answers a small set of text queries derived from multiple task-generic prompts $p_{i}^{g}$ and returns foreground and background category tags, noted as $tag^{fg}$ and $tag^{bg}$. (2) Instance Mask Generator. Grounding DINO converts each foreground tag into multiple instance-aware bounding boxes, while the proposed Single-Foreground Multi-Background Prompting (SFMBP) strategy samples region-constrained foreground and multi-background points inside every box from CLIP-derived heatmaps; all box-and-points pair forms multiple instance-level visual prompts for SAM, which produces a candidate instance mask. (3) Self-consistency Instance Mask Voting strategy. Among all candidate instance masks, the one closest to their mean is selected as the final COS result. With all models kept frozen, the IAPF delivers instance-aware, zero-shot COS at negligible additional cost.
Text Prompt Generator
The Text Prompt Generator leverages the advanced visual question capabilities of the Multimodal Large Language Models (MLLM) (Liu et al. 2024) to convert a task-generic prompt $p_{i}^{g}$ to text prompts tailored for each input image. Formally, given an input image $X\in\mathbb{R}^{H\times W\times 3}$ representing the visual content, a textual query $Q$ is fed into the MLLM with parameters $\theta$ to generate a textual response sequence $y$. The generated response $y$ explicitly characterizes details relevant to the query. The generation process follows an auto-regressive probabilistic modeling framework, sequentially predicting tokens based on the visual context provided by $X$, the textual query $Q$, and previously generated tokens. Mathematically, the entire response generation can be represented by the following joint probability:
$p(y\mid X,Q;\theta)$ $=\prod_{t=1}^{T}p(y_{t}\mid y_{<t},X,Q;\theta),$ (1)
where $y_{t}$ denotes the token generated at time step $t$, $y_{<t}$ represents the sequence of tokens generated before time step $t$, and $T$ is the total length of the generated response.
For simplicity, we adopt the same protocol of GenSAM (Hu et al. 2024a) for extracting the text prompts required by the Instance Mask Generator. Concretely, for each task-generic prompt $p^{g}_{i}$ we construct two textual queries, $Q_{i}^{f}$ and $Q_{i}^{b}$, and submit them to a frozen MLLM:
$tag_{i}^{fg}=$ $\text{MLLM}(X,C_{i},Q^{f}_{i}),$ (2)
$tag_{i}^{bg}=$ $\text{MLLM}(X,C_{i},Q^{f}_{i},tag_{i}^{fg},Q^{b}_{i}),$ (3)
where $C_{i}$ is the image caption first produced by the MLLM. The two textual queries are instantiated as $Q_{i}^{f}\!:$ “Name of the $p^{g}_{i}$ in one word.” and $Q_{i}^{b}\!:$ “Name of the environment of the $p^{g}_{i}$ in one word.”. The single-word answers returned by the model are recorded as the foreground tag $tag_{i}^{fg}$ and background tag $tag_{i}^{bg}$, respectively.
Instance Mask Generator
Methods Venue Setting NC4K COD10K-TEST CAMO-TEST
$S_{\alpha}$ $F_{\beta}^{\omega}$ $M$ $E_{m}^{\phi}$ $S_{\alpha}$ $F_{\beta}^{\omega}$ $M$ $E_{m}^{\phi}$ $S_{\alpha}$ $F_{\beta}^{\omega}$ $M$ $E_{m}^{\phi}$
UCOS-DA (Zhang and Wu 2023) ICCV23 $U$ .755 .656 .085 .819 .689 .513 .086 .739 .700 .605 .127 .784
SdalsNet (Shou et al. 2025) AAAI25 $U$ .739 .642 .085 .824 .697 .525 .072 .780 .697 .601 .117 .799
EASE (Du et al. 2025) CVPR25 $U$ .800 .735 .056 .884 .773 .656 .040 .866 .749 .684 .098 .831
WS-SAM (He et al. 2023a) NeurIPS23 $P$ .813 .734 .057 .859 .790 .663 .039 .856 .718 .602 .102 .757
P-COD (Chen et al. 2024a) ECCV24 $P$ .822 .748 .051 .889 .784 .650 .042 .859 .798 .727 .074 .872
SAM-COD (Chen et al. 2024b) ECCV24 $P$ .858 .802 .041 .918 .831 .725 .031 .901 .820 .760 .066 .885
BoxSAM (Li et al. 2025a) IVC25 $P$ .810 - .069 .823 .808 - .042 .851 .745 - .102 .749
SS (Zhang et al. 2020) CVPR20 $S$ .723 .584 .090 .803 .684 .466 .068 .774 .684 .538 .119 .760
SCWS (Yu et al. 2021) AAAI21 $S$ .757 .659 .072 .846 .708 .535 .059 .813 .719 .616 .105 .810
CRNet (He et al. 2023b) AAAI23 $S$ .775 .688 .063 .855 .733 .576 .049 .832 .735 .641 .092 .815
WS-SAM (He et al. 2023a) NeurIPS23 $S$ .829 .757 .052 .886 .803 .680 .038 .877 .759 .667 .092 .818
MINet (Niu et al. 2024) ACM MM24 $S$ .793 .709 .061 .869 .749 .596 .049 .840 .750 .669 .091 .840
WSMD (Zha et al. 2024) AAAI24 $S$ .797 .700 .064 .868 .761 .600 .049 .839 .793 .704 .079 .866
SAM-COD (Chen et al. 2024b) ECCV24 $S$ .859 .795 .039 .912 .833 .728 .029 .904 .836 .779 .060 .903
BoxSAM (Li et al. 2025a) IVC25 $S$ .860 - .052 .894 .836 - .044 .883 .830 - .084 .860
SAM-COD (Chen et al. 2024b) ECCV24 $B$ .867 .813 .037 .923 .842 .745 .028 .914 .837 .786 .062 .901
BoxSAM (Li et al. 2025a) IVC25 $B$ .877 - .037 .925 .857 - .027 .919 .859 - .057 .908
LLaVA1.5+SAM (Liu et al. 2024) CVPR23 ZSZS .609 .543 .253 .676 .653 .551 .193 .719 .571 .526 .292 .631
CLIP_Surgey+SAM (Li et al. 2025b) PR25 ZSZS .712 .626 .152 .755 .661 .531 .174 .699 .671 .601 .179 .729
Grounded SAM (Ren et al. 2024) Arxiv24 ZSZS .771 .721 .135 .806 .745 .667 .146 .784 .683 .654 .203 .723
MMCPF (Tang et al. 2024) ACM MM24 ZSZS .768 .681 .082 - .733 .592 .065 - .749 .680 .100 -
GenSAM (Hu et al. 2024a) AAAI24 ZSZS .805 .745 .066 .863 .783 .682 .055 .845 .727 .648 .105 .799
ProMaC (Hu et al. 2024b) NeurIPS24 ZSZS .814 .762 .056 .884 .803 .710 .042 .875 .745 .685 .100 .830
RDVP-MSD (Yin et al. 2025) ACM MM25 ZSZS .823 .764 .056 .873 .825 .743 .038 .877 .796 .750 .081 .848
ours ZSZS .862 .828 .043 .916 .856 .799 .033 .917 .803 .768 .081 .864
Table 1: Quantitative comparison of COS methods across three benchmark datasets. The table compares state-of-the-art methods based on unsupervised, weakly-supervised, and training-free setups. ’$U$’ indicates unsupervised settings, while the weakly-supervised settings include ’$P$’ (point), ’$S$’ (scribble), and ’$B$’ (box). ’ZSZS’ represents zero-shot training-free approaches applied to COS. ’-’ denotes the absence of explicit mention and unavailable code in the corresponding paper. ’↑’ indicates higher is better, and ’↓’ indicates lower is better. The best results of different settings are highlighted in bold.
Existing training-free COS pipelines (Yin et al. 2025; Hu et al. 2024a, b; Tang et al. 2024) predominantly rely on a single semantic-level box prompt per image, which limits SAM (Kirillov et al. 2023) to return only one coarse mask, thereby hindering comprehensive segmentation when multiple camouflaged instances coexist. To overcome this bottleneck, the proposed Instance Mask Generator serves as the crucial bridge between the text prompts derived from the Text Prompt Generator and the segmentation masks output by the SAM. The process involves three frozen steps: (1) Box prompting: Grounding DINO (Liu et al. 2023) converts each foreground tag into multiple instance-level bounding box prompts. (2) Point prompting: Within every box prompt, the proposed Single-Foreground Multi-Background Prompting (SFMBP) strategy samples high-confidence foreground points and equally confident multi-background points, yielding instance-level point prompts. (3) Mask Prediction: Each instance-level visual prompt (i.e., a box–points pair) is fed to SAM for producing a mask for an instance; aggregating all masks yields a candidate instance mask that covers all camouflaged instances in the image.
Box Prompting
Existing training-free COS methods often fail to segment objects in scenarios containing multiple discrete camouflaged instances, as they typically provide a single bounding box per image for SAM to segment. According to SAM, accurate instance segmentation requires explicit object bounding box prompts (Kirillov et al. 2023). Therefore, the simplest solution to this problem is to provide an instance-level bounding box prompt for each camouflaged instance in the image. To this end, we introduce Grounding DINO (Liu et al. 2023) and leverage its instance-aware detection capabilities. Specifically, for each foreground tag $tag_{i}^{fg}$, we apply Grounding DINO to the pair $(X,tag_{i}^{fg})$ to generate multiple instance-level bounding box prompts $\mathcal{B}_{i}$:
$\mathcal{B}_{i}=\text{G-DINO}(X,tag_{i}^{fg}),$ (4)
where $\mathcal{B}_{i}$ represents the $N$ instance-level bounding box prompts generated by Grounding DINO for the $i$-th foreground tag $tag_{i}^{fg}$. This approach replaces the ”single box per image” assumption with detector-driven coverage of all camouflaged instances.
Point Prompting
RDVP-MSD (Yin et al. 2025) introduced the Region-Constrained (RC) strategy, where point prompts are sampled within a single box per image, thereby alleviating the issue that previous approaches (Hu et al. 2024a, b) mostly relied on image-level heatmaps to generate point prompts. However, this approach, like others, is still limited by two key factors: (1) the camouflage environment is often complex and diverse, making it difficult to generate sufficient background point prompts using a single background tag; (2) the bounding box is typically a single box per image, severely limiting its efficacy in scenarios containing multiple discrete camouflaged instances.
To mitigate these issues, we propose Single-Foreground Multi-Background Prompting (SFMBP) strategy, which leverages multi-background cues along with instance-level bounding box prompts to sample foreground and background points more effectively. Specifically, Spatial CLIP (Hu et al. 2024a) is used to generate a foreground heatmap $H^{f_{i}}$ and multiple background heatmaps $H^{b}=\{H^{b_{1}},H^{b_{2}},\dots,H^{b_{L}}\}$, where each $H^{b_{l}}$ corresponds to a specific background tag $tag^{b_{l}}$ from $\text{set}(tag^{bg})$, and $L$ is the length of $\text{set}(tag^{bg})$:
$H^{f_{i}}$ $=\text{Spatial CLIP}(X,tag_{i}^{fg}),$ (5)
$H^{b_{l}}$ $=\text{Spatial CLIP}(X,tag^{b_{l}}),\quad l=1,2,\dots,L,$ (6)
where $H^{f_{i}}$ is the foreground heatmap generated for the foreground tag $tag_{i}^{fg}$, and each $H^{b_{l}}$ is a background heatmap corresponding to a different background tag $tag^{b_{l}}\in\text{set}(tag^{bg})$.
These heatmaps encode pixel-level confidence scores for the foreground and background categories. By adaptively sampling high-confidence points from both the $i$-th foreground heatmap and all background heatmaps inside every instance-level bounding box $\mathcal{B}_{i,n}\in\mathcal{B}_{i}$, SFMBP ensures the acquisition of accurate instance-level point prompts, addressing both the issue of the complexity of diverse background environments and the single box limitation. Formally, the point sampling process can be described as follows:
$p_{i,n}^{fg}$ $=\left\{(x,y)\mid Norm(H^{f_{i}}\mid\mathcal{B}_{i,n})_{(x,y)}\geq 0.9\right\},$ (7)
$p_{i,n}^{\mathrm{bg}}$ $=\bigcup_{l=1}^{L}\left\{(x,y)\mid Norm(H^{b_{l}}\mid\mathcal{B}_{i,n})_{(x,y)}\geq 0.9\right\},$ (8)
where $p_{i,n}^{fg}$ represents the foreground point prompts inside the $n$-th box $B_{i,n}$ with high confidence, and $p_{i,n}^{\mathrm{bg}}$ represents the background point prompts inside the same box, sampled from the multiple background heatmaps. The $Norm(\cdot)$ denotes a normalization function, ensuring pixel-level confidence scores are normalized within a standard range(i.e., $[0,1]$).
Mask Prediction
[…]
$m_{i,n}=\text{SAM}(X\mid\langle\mathcal{B}_{i,n},\,p^{fg}_{i,n},\,p^{bg}_{i,n}\rangle),n=1,\dots,N,$ (9)
where the $N$ denotes the number of instance-level bounding boxes corresponding to the $tag_{i}^{fg}$. The resulting masks $\{m_{i,1},\ldots,m_{i,N}\}\in\{0,1\}^{H\times W}$ delineate all camouflaged instances associated with $tag^{fg}_{i}$. To yield a candidate instance mask that faithfully covers every instance, we stack them along a new dimension:
$mask_{i}=[m_{i,1};m_{i,2};...;m_{i,N}],$ (10)
where the $mask_{i}\in\{0,1\}^{N\times H\times W}$.
Self-consistency Instance Mask Voting
Figure 3: Qualitative comparison of the proposed IAPF with three main training-free COS methods. From left to right, as the number of camouflaged instances in the scene increases, existing training-free methods fail to segment all objects. In contrast, the IAPF consistently produces high-quality instance masks, even when dealing with multiple camouflaged instances.
Methods Venue Setting NC4K COD10K-TEST
APAP $AP_{50}$ $_{75} APAP$ $AP_{50}$ $AP_{75}$
OSFormer (Pei et al. 2022) ECCV22 $F$ .444 .737 .451 .420 .713 .428
CLPNet (Cui, Sun, and Yuan 2023) ICIP23 $F$ .451 .748 .455 .433 .719 .432
QFormer (Dong et al. 2023) ACM MM23 $F$ .501 .768 .528 .454 .718 .479
DCNet (Luo et al. 2023) CVPR23 $F$ .528 .771 .565 .453 .707 .475
AQSFormer (Dong et al. 2024) ACM MM24 $F$ .505 .768 .535 .465 .738 .485
Mask2Camouflage (Phung and Shuai 2024) ACCV24 $F$ .538 .776 .583 .468 .725 .490
TPNet (He et al. 2024) ACM MM24 II .214 .483 .166 .183 .418 .143
Ours ZSZS .511 .807 .555 .448 .748 .475
Table 2: Quantitative evaluation on CIS across NC4K and COD10K-TEST datasets. ’F’ and ’I’ denote fully-supervised and image-level supervised methods, respectively. ‘ZS’ denotes zero-shot training-free setting.
Following previous methods (Yin et al. 2025), we also adopt a self-consistency approach to select the most robust segmentation result. Specifically, we introduce a Self-consistency Instance Mask Voting (SIMV) mechanism to select the best instance segmentation mask rather than a semantic mask. In the SIMV mechanism, we need a semantic mask that is directly comparable across repeated runs. To achieve this, we convert the instance mask $mask_{i}$ into a semantic mask $M_{i}\in\{0,1\}^{H\times W}$ by performing a per-pixel logical OR $\bigvee$ over the instance dimension $N$:
$M_{i}(x,y)=\bigvee_{n=1}^{N}mask_{i}(n,x,y).$ (11)
We repeat the pipeline $I$ times (default $3$) with synonymic task-generic prompts $p_{i}^{g}$, resulting in a set of candidate semantic masks $\{M_{1},\dots,M_{I}\}$. To quantify their consistency, we apply the self-consistency rule:
$\hat{i}=\arg\min_{i}\;\left|M_{i}-\frac{1}{I}\sum_{i=1}^{I}M_{i}\right|.$ (12)
The output $M_{\hat{i}}$ and the corresponding $mask_{\hat{i}}$ represent the final COS and CIS predictions, respectively.
Experiments
Experiment Settings
Datasets and Evaluation Metrics
For COS, we evaluate on widely-used datasets, including NC4K ( (Lv et al. 2021), containing 4,121 images), COD10K-TEST ( (Fan et al. 2020), with 2,026 images), and CAMO-TEST ( (Le et al. 2019), consisting of 250 images). We use S-measure ($S_{\alpha}$(Fan et al. 2017), weighted F-measure ($F^{\omega}_{\beta}$), mean absolute error ($M$), and mean E-measure ($E_{m}^{\phi}$) to evaluate the performance. High-performing COS methods generally achieve higher values of $S_{\alpha}$, $F^{\omega}_{\beta}$, and $E_{m}^{\phi}$, along with a lower $M$ score. For CIS, we evaluate on NC4K (Lv et al. 2021) and COD10K-TEST (Fan et al. 2020) and report Average Precision ($AP$), $AP_{50}$, and $AP_{75}$. Higher values of $AP$, $AP_{50}$, and $AP_{50}$ indicate superior camouflaged instance segmentation performance.
Implementation details
In our experiments, we employ LLaVA-1.5-13B (Liu et al. 2024) as the MLLM. For the CLIP model, we use the CLIP (Radford et al. 2021) from the CS-ViT-L/14@336px variant. For the SAM, we deploy the HQ-SAM (Ke et al. 2023) built on the ViT-H version. Our method is entirely train-free, requiring no fine-tuning. All experiments are conducted using the PyTorch framework, and we utilize two NVIDIA GeForce RTX 3090 GPUs with 24 GB of memory each for evaluation. The reported results are based on a single execution without repeated trials or statistical variation analysis.
Comparison with State-of-the-Art Methods
Quantitative and Qualitative Comparison
We compare the proposed IAPF with unsupervised, weakly-supervised, and training-free methods on the NC4K, COD10K-TEST, and CAMO-TEST datasets, as shown in Table 1. In the zero-shot training-free setting, our method achieves significant improvements over all compared methods, including higher $S_{\alpha}$, $F_{\beta}^{\omega}$, and $E_{m}^{\phi}$ values. This performance improvement is mainly due to the proposed stance Mask Generator, which generates multiple instance-level bounding boxes, and the SFMBP strategy. These innovations enable more accurate segmentation, especially in challenging scenarios with multiple camouflaged instances. Compared to unsupervised and weakly-supervised methods, the proposed IAPF achieves competitive performance without any form of supervision or training.
Qualitative results in Figure 3 show that, as the number of camouflaged instances increases, previous training-free COS methods struggle to segment the objects accurately, while our method maintains strong performance across all scenarios.
Evaluation on Camouflaged Instance Segmentation
We further evaluate our method on the Camouflaged Instance Segmentation (CIS) task, reporting quantitative results on the NC4K and COD10K-TEST datasets in Table 2. Compared to recent fully-supervised (’$F$’) and image-level supervised (’$I$’) CIS approaches, our method achieves competitive or even superior performance, notably attaining the highest $AP_{50}$ scores on both NC4K and COD10K-TEST datasets. Remarkably, our results are obtained entirely in a training-free manner using only a single task-generic prompt, underscoring the effectiveness of the instance-aware prompting strategy for accurate instance segmentation without reliance on costly annotations.
Method’s Variants NC4K COD10K-TEST
$S_{\alpha}$ $F_{\beta}^{\omega}$ $M$ $E_{m}^{\phi}$ $S_{\alpha}$ $F_{\beta}^{\omega}$ $M$ $E_{m}^{\phi}$
(1) wo TPG .855 .816 .046 .909 .840 .768 .036 .902
(2.1) wo G&wo SFMBP .788 .707 .074 .840 .760 .626 .064 .815
(2.2) wo G&w SFMBP .797 .714 .062 .850 .766 .627 .053 .822
(2.3) w G&wo SFMBP .847 .802 .051 .900 .835 .760 .041 .894
(3) wo SIMV .850 .812 .042 .902 .844 .780 .031 .904
(4) Ours .862 .828 .043 .916 .856 .799 .033 .917
Table 3: Ablation study on the effectiveness of the IAPF.
Repeat NC4K COD10K-TEST
$S_{\alpha}$ $F_{\beta}^{\omega}$ $M$ $E_{m}^{\phi}$ $S_{\alpha}$ $F_{\beta}^{\omega}$ $M$ $E_{m}^{\phi}$
1 .850 .812 .042 .902 .844 .780 .031 .904
2 .857 .820 .045 .912 .852 .789 .034 .913
3 .862 .828 .043 .916 .856 .799 .033 .917
4 .861 .827 .044 .914 .855 .798 .033 .915
5 .859 .824 .044 .912 .855 .796 .034 .914
Table 4: Ablation study on the effectiveness of repetition number II.
Methods Ins COD10K-TEST
$S_{\alpha}$ $F_{\beta}^{\omega}$ $M$ $E_{m}^{\phi}$
(1) .789 .690 .055 .852
GenSAM (2) .730 -7.48% .622 -9.86% .064 -16.36% .820 -3.76%
(3) .632 -19.90% .443 -35.80% .077 -40.00% .700 -17.84%
(1) .814 .725 .040 .881
ProMaC (2) .760 -6.63% .667 -8.00% .051 -27.50% .855 -2.95%
(3) .634 -22.11% .457 36.97% .073 -82.50% .750 -14.87%
(1) .833 .756 .038 .886
RDVP-MSD (2) .774 -7.08% .674 -10.85% .042 -10.53% .834 -5.87%
(3) .648 -22.21% .455 -39.81% .079 -107.89% .672 -24.15%
(1) .860 .804 .033 .918
Ours (2) .840 -2.33% .788 -1.99% .032 +3.03% .921 +0.33%
(3) .777 -9.65% .697 -13.31% .050 -51.52% .846 -7.84%
Table 5: Performance comparison of different camouflaged instance counts on COD10K-TEST.
Ablation Study
Effectiveness of the Modules
We perform ablation studies to examine the effectiveness of each key component in the IAPF framework: Text Prompt Generator (TPG), Grounding DINO (G), Single-Foreground Multi-Background Prompting (SFMBP), and Self-consistency Instance Mask Voting (SIMV). As shown in Table 3, removing TPG (setting 1), G and SFMBP (settings 2.1-2.3), or SIMV (setting 3) clearly degrades performance, verifying the necessity of each module. Moreover, comparisons between settings (2.1 vs. 2.3) and (2.2 vs. 4) further demonstrate that introducing Grounding DINO not only grants the IAPF the critical instance-awareness capability but also significantly improves overall segmentation accuracy by allowing the model to localize more camouflaged instances. The complete method (setting 4) consistently outperforms all variants on both NC4K and COD10K-TEST datasets, confirming that each proposed component contributes distinctly and effectively to the final COS performance.
Effectiveness of Repetition Number
We evaluate the impact of the repetition number $I$ on the stability and performance of our method. As shown in Table 4, increasing the number of repetitions improves the segmentation results. The best performance is achieved when $I=3$, with higher repetitions providing diminishing returns.
Performance Comparison of Different Camouflaged Instance Counts
Table 5 reports COS performance on COD10K-TEST after stratifying images by the number of camouflaged instances: (1) single, (2) double, and (3) three-or-more objects. For each metric, we also list the relative change $(\%)$ w.r.t. the single-instance baseline (row 1 of every method).
The proposed IAPF maintains consistently high accuracy as instance density grows. The larger percentage swing in $M$ for all methods is expected because the absolute MAE values are very small; nonetheless, the absolute MAE of the IAPF remains the lowest. These results confirm that the instance-aware prompting pipeline––particularly the Grounding DINO–based box prompts and SFMBP point sampling––provides strong resilience to cluttered scenes with many camouflaged objects.
Conclusion
We introduced the IAPF, the first training-free framework that converts a single task-generic prompt into instance-level guidance for SAM, enabling zero-shot training-free segmentation of multiple camouflaged objects. By coupling Grounding DINO-based instance-level box prompts and the SFMBP point-sampling strategy, the IAPF delivers fine-grained COS without any additional training or annotations. Extensive experiments on standard COS benchmarks show that our method consistently outperforms prior unsupervised, weakly-supervised, and training-free COS approaches.